Exercise 15-22 Pina Companys ledger shows the following balances on December 31, 2017 4% Preferred Stock-$10 par value, outstanding 18,000 shares Common Stock-$100 par value, outstanding 27,500 shares Retained Earnings $180,000 2,750,000 577,000 Assuming that the directors decide to declare total dividends in the amount of $380,000, determine how much each class of stock should receive under each of the conditions stated below. One years dividends are in arrears on the preferred stock. (a) The preferred stock is cumulative and fully participating. (Round the rate of participation to 4 decimal places, e.g. 1.4278%. Round answers to 0 decimal places, e.g $38,487.) Preferred Common

(b) The preferred stock is noncumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.) Preferred Common (c) The preferred stock is noncumulative and is participating in distributions in excess of a 6% dividend rate on the common stock. (Round the rate of participation to 4 decimal places, eg. 1.4278%. Round answers to 0 decimal places, eg. $38,487.) Preferred Common
Answer #1
Given that-
4% preferred stock , 18000 outstanding 180000
Common stock, 27500 outstanding 2750000
retained earning 577000
Answer a)
Arrear dividend for preferred stock = 7200 7200
Current year dividend for preferred stock = 7200 7200
Total dividend to preferred stock = 14400
Dividend to common sharestock = 2750000*4% 110000
Total distributable amount = 380000
Less : preferred and common dividend = 124400 124400
Remaining amount to be distributed proportionately = 255600
Share of preferred stock = 255600*180000/(180000+2750000) 15702
Share of common stock = 255600*2750000/(180000+2750000) 239898
Total 255600
Total dividend paid under a)
Preferred 30102 30102
Common 349898 349898
Answer b)
In this situation preferred stock will not get any arrear share and not excess dividend
Current year dividend for preferred stock = 7200 7200
Common dividend = 380000-7200 372800
Preferred 7200
Common 372800
Answer c)
Current year dividend for preferred stock = 7200 7200
Total dividend to preferred stock = 7200
Dividend to common sharestock = 2750000*6% 165000
Total distributable amount = 380000
Less : preferred and common dividend = 172200 172200
Remaining amount to be distributed proportionately = 207800
Share of preferred stock = 207800*180000/(180000+2750000) 12765.9
Share of common stock = 207800*2750000/(180000+2750000) 195034
Total 207800
Total dividend paid under c)
Preferred 19965.9
Common 360034